OpenStudy (anonymous):

Find the coordinates of the x-intercept. -x - 2y = 6

OpenStudy (anonymous):

Hmm.. there's probably a better way to do this, but here's how I'd solve it: First find the y intercept by converting to the y intercept form of a line. -x - 2y = 6 -2y=x+6 2y=-x-6 y=-1/2x-3 So we know that the y-intercept is -3. Now to go up once, we go over 2 to the left. So to go up 3, (to find the x-intercept) we need to go over 6 to the left. So our x intercept is -6.

OpenStudy (anonymous):

Sorry. you want the coordinates of the x intercept. It would just be (-6, 0) because where it intercepts the x axis y will equal 0.

OpenStudy (ipwnbunnies):

Another, simpler, way to solve it is the set y = 0.

OpenStudy (anonymous):

LOL! Thanks iPwnBunnies. I didn't even think of that.